OpenStudy (anonymous):

Is (-1,-5) the solution to 6x-5<= 11x + y ? If o, how do I graph it properly?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

no its not the solution to it 6x-5<= 11x + y just plug the x=-1, y=-5 6(-1)-5<= 11(-1) + -5 -6 -5 <= -11-5 -11 <=-16 false....bec -16< -11

OpenStudy (anonymous):

Try to see -5 in terms of x and y first... 6x-5<= 11x + y -5<= 11x + y -6y -5<= 5x + y If (-1, -5) is a solution then bu replacingx with -1 and y with -5 we get... -5<5(-1) + (-5) which yields ... -5<-5 -5 which simplifiest to ... -5 <-10 This you will see is an untrue statement and as such the coordinate paire do not satisfy this particular inequality.

OpenStudy (anonymous):

How to graph it properly ... Write y in terms of x and then feed 3 values in to the x variable. 6x-5<= 11x + y impliest that ... 6x -11x -5 <= y or even better ... -5x-5 <= y If x = 0 say, then y = -5 hence (0, -5) is on the line If x = -2 say, then y = 5 hence (-2, 5) is on the line If x = -3 say, then y = 10 hence (-3, 10) is on the line Now check the attached jpeg to see the graph The shaded area represents all coordinates that make the inequality true THE BROKEN GRAPH LINE SHOWS THAT THE COORDINATES ON THAT LINE ARE NOT PERMISSABLE.